OpenStudy (anonymous):

Select one of the factors of x2y3 - 11x2y + 6y2 - 66

OpenStudy (raden):

x^2y^3 - 11x^2y + 6y^2 - 66 = x^2y(y^2-11) + 6(y^2-11) = (y^2-11)(x^2y+6)